find the period, range, and amplitude of the cosine function y=(3)/(2)cos(t/2)
Answer by lwsshak3(11628)

find the period, range, and amplitude of the cosine function y=(3)/(2)cos(t/2)
..
y=(3)/(2)cos(t/2)
Standard form for cosine function:
y=Acos(Bx-C), A=amplitude, Period=2π/B, C/B=phase-shift
For given cosine function:
Amplitude=3/2
Period=2π/B=2π/(1/2)=4π
Range=[-3/2,3/2]
Phase-shift=0